Capital accumulation: Kt+1 −Kt =∆Kt+1 =It −δ 1) Please discuss the intuition of the capital accumulation...

  • Capital accumulation: Kt+1 −Kt =∆Kt+1 =It −δ

1) Please discuss the intuition of the capital accumulation equation above.

2) Please derive the equation that describes the evolution of capital stock per capita
Answer #1

1)

Capital accumulation refers to a change in the availability of capital. It can be measured on a yearly basis. It is a net addition to the capital stock.

When the new investment is made, it replaces depreciation or investment above the depreciation is called capital accumulation. If the investment is less than depreciation, there will be a decline in the capital but if investment tends to be larger than the depreciation, there will be an accumulation of capital.

2)

Kt+1 −Kt =∆Kt+1 =It −δ

or

I = s(y)

∆K = s(Y) - dK ( d = Depreciation)

or

Y = f(k)

Thus,

∆K = s(k) - dK.

This is called per capita rise in the capital stock.